# PATCH! Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Proficiency in 8th Grade Mathematics

### v2

@@ Abstract
Many existing benchmarks of large (multimodal) language models (LLMs) focus on measuring LLMs’ academic proficiency, often with also an interest in comparing model performance with human test takers. While these benchmarks have proven key to the development of LLMs, they suffer from several limitations, including questionable measurement quality (e.g., Do they measure what they are supposed to in a reliable way?), lack of quality assessment on the item level (e.g., Are some items more important or difficult than others?) and unclear human population reference (e.g., To whom can the model be compared?). In response to these challenges, we propose leveraging knowledge from psychometrics - a field dedicated to the measurement of latent variables like academic proficiency - into LLM benchmarking. We make three primary contributions. First, we introduce PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs. PATCH addresses the aforementioned limitations, presenting a new direction for LLM benchmark research. Second, we implement PATCH by measuring GPT-4 and Gemini-Pro-Vision’s proficiency in 8th grade mathematics against 56 human populations. We show that adopting a psychometrics-based approach yields evaluation outcomes that diverge from those based on existing benchmarking practices. Third, we release 4 high-quality datasets to support measuring and comparing LLM proficiency in grade school mathematics and science against human populations.
@@ 1 Introduction
Large language models (LLMs), including their multimodal variants like vision language models, have witnessed significant advancements in recent years. These models are typically evaluated on established benchmarks that assess their performance across a diverse set of tasks, including commonsense reasoning (e.g., HellaSwag by Zellers et al. (2019), WinoGrande by Sakaguchi et al. (2021)), coding (HumanEval by Chen et al. (2021), Natural2Code by Google (2023), and academic proficiency. Academic proficiency, in particular, has become a crucial part of LLM evaluation, as evidenced by the large number of related benchmarks (e.g., MMLU by Hendrycks et al. (2021), ARC by Clark et al. (2018), GSM8K by Cobbe et al. (2021), DROP by Dua et al. (2019), MATH by Hendrycks et al. (2021)), and recent model technical reports’ focus on them (e.g., OpenAI, 2023; Google, 2023). In these benchmarks, LLM performance is also often contrasted with human performance.
@@
Despite the success of existing benchmarks in advancing LLM research, they are not without limitations. The first concern is measurement quality: Do these benchmarks measure what they are supposed to in a reliable way? Many benchmarks are created via crowd-sourced knowledge, by asking a convenience group of individuals (e.g., crowd workers, paper authors) to create new test items (e.g., GSM8K, DROP) or collecting them from (often undocumented) sources (e.g., websites, textbooks, school exams) (e.g., MATH, MMLU, ARC). Without domain expert input and rigorous testing of item quality, undesirable outcomes can occur, including a mismatch between a benchmark and its claimed measurement goal, missing information in a question, wrong answer keys, and low data annotation agreement (Nie et al., 2020). [^1] We avoid calling out specific datasets here, but a quick Internet search would reveal many blogs reporting large percentages of errors in existing LLM benchmarks.
@@
Second, current benchmarks do not account for differences across test items, such as item discriminativeness [^2] In psychometrics, the term “item discrimination” is used. However, given the ambiguity and negative connotation of “discrimination”, we adopt “discriminativeness” instead. and difficulty (see Section 3.1). For example, consider three items A (easy), B (hard) and C (hard). While answering correctly to A and B would result in the same accuracy score as answering correctly to B and C, the latter (i.e., answering correctly to more difficult items) would imply higher proficiency. Furthermore, items that are too easy or too difficult (i.e., low discriminativeness) will fail to differentiate models of different proficiency levels. Thus, without accounting for item differences, benchmarking results, especially model rankings, can be misleading.
@@
Third, while many benchmarks are used to compare LLMs against humans, the human population to be compared is unclear (Tedeschi et al., 2023). For instance, human performance in MATH is based on the paper’s authors; in MMLU, crowd workers; in MATH, 6 university students. Using such convenience samples (with none to little information about sample characteristics), the resulting human performance is local to that specific sample and cannot be generalised to other human samples or specific populations.
@@
To address these challenges, we propose integrating insights from psychometrics - a field dedicated to the measurement of latent variables like cognitive abilities and academic proficiency - into LLM benchmarking processes. In particular, we draw on two research areas in psychometrics: item response theory (see Section 3.1) and test development (see Section 3.2 and 3.3). The former can help to estimate academic proficiency more accurately than common practice in LLM benchmarks (e.g., means, percentages, accuracy scores). It can also provide diagnostic information about the quality of each test item. The latter, test development knowledge, can help to build high quality LLM benchmarks where comparison to specific human populations can be made.
@@
Our paper makes three primary contributions. First, we present PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs, which addresses the aforementioned limitations of existing benchmarks. Second, we demonstrate the implementation of PATCH by testing GPT-4 and Gemini’s proficiency in 8th grade mathematics using the released test items and data from Trends in International Mathematics and Science Study [^3] http://timssandpirls.bc.edu/timss2015/encyclopedia/ (TIMSS) 2011. We show empirically that a psychometrics-based approach can lead to evaluation outcomes that diverge from those obtained through conventional benchmarking practices and are more informative, underscoring the potential of psychometrics to reshape the LLM benchmarking landscape. Third, we make our benchmark dataset and evaluation code [^4] Available at https://github.com/fqixiang/patch_llm_benchmarking_with_psychometrics based on TIMSS 2011 available to future researchers, along with three other mathematics and science datasets based on TIMSS 2011 and 2008555Available at https://zenodo.org/records/12531906.
@@ 2 Related Work
We are not the first to propose leveraging psychometrics for research on LLMs and other areas in NLP. For instance, psychometric scales have been used to examine the psychological profiles of LLMs such as personality traits and motivations (Huang et al., 2024; Pellert et al., 2023; Dillion et al., 2023). The text in these scales can also be used to improve encoding and prediction of social science constructs like personality traits (Kreuter et al., 2022; Vu et al., 2020; Yang et al., 2021; Fang et al., 2023a). Psychometrics-based reliability and validity tests have also been proposed or/and used to assess the quality of NLP bias measures (Du et al., 2021; van der Wal et al., 2024), text embeddings (Fang et al., 2022), political stance detection (Sen et al., 2020), annotations (Amidei et al., 2020), user representations (Fang et al., 2023b), and general social science constructs (Birkenmaier et al., 2023).
The most closely related work to our paper is the use of IRT models in NLP for constructing more informative test datasets (Lalor et al., 2016), comparison of existing evaluation datasets and instances (e.g., difficulty, discriminativeness) (Sedoc & Ungar, 2020; Vania et al., 2021; Rodriguez et al., 2021; Lalor et al., 2018; Rodriguez et al., 2022), as well as identification of difficult instances from training dynamics (Lalor & Yu, 2020; Lalor et al., 2019). Our work distinguishes itself from these papers in two aspects. First, we do not apply IRT to existing LLM datasets/benchmarks. Instead, we introduce a framework for benchmarking LLMs by leveraging both IRT and test development knowledge from psychometrics. The goal of this framework is to generate new, high-quality benchmarks for LLMs that warrant valid comparison with human populations. Second, we demonstrate our framework with a mathematics proficiency test validated on 56 human populations, and compare LLM performance with human performance. To the best our knowledge, we are the first to apply psychometrically validated (mathematics) proficiency tests to LLMs and make valid model versus human comparison.
@@ 3.1 Item Response Theory
Item response theory (IRT) refers to a family of mathematical models that describe the functional relationship between responses to a test item, the test item’s characteristics (e.g., item difficulty and discriminativeness) and test taker’s standing on the latent construct being measured (e.g., proficiency) (AERA et al., 2014). Unlike classical test theory and current LLM benchmarks, which focus on the total or mean score of a test, IRT models takes into account the characteristics of both the items and the individuals being assessed, offering advantages like more accurate estimation of test takers’ proficiency, and item quality diagnostics. As such, IRT models have gained widespread adoption in various fields, including education, psychology, and healthcare, where precise measurement and assessment are crucial.
@@
$P\left(x_{i}=1\mid\theta,a_{i},b_{i},c_{i}\right)=c_{i}+\frac{1-c_{i}}{1+\exp\left(-1.7\cdot a_{i}\cdot\left(\theta-b_{i}\right)\right)}\equiv P_{i,1}\left(\theta\right)$ (1)
where $x_{i}$ is the scored response to item $i$ ($1$ if correct and 00 if incorrect); $\theta$ is the proficiency of the test taker, where higher proficiency has a greater probability of responding correctly; $a_{i}$ is the slope parameter of item $i$, characterising its discriminativeness (i.e., how well the item can tell test takers with higher $\theta$ from those with lower $\theta$) [^6] The number $1.7$ is a scaling parameter to preserve historical interpretation of parameter $a_{i}$ on the normal ogive scale (Camilli, 1994). Also applies to 2PL and GPC models.; $b_{i}$ is the location parameter of item $i$, characterising its difficulty; $c_{i}$ is the lower asymptote parameter of item $i$, reflecting the chances of test takers with very low proficiency selecting the correct answer (i.e., guessing). Correspondingly, the probability of an incorrect response to item $i$ is: $P_{i,0}=P\left(x_{i}=0\mid\theta_{k},a_{i},b_{i},c_{i}\right)=1-P_{i,1}\left(\theta_{k}\right)$. The 2PL model has the same form as the 3PL model (Equation 1), except that the $c_{i}$ parameter is fixed at zero (i.e., no guessing).
@@
$P\left(x_{i}=l\mid\theta,a_{i},b_{i},d_{i,1},\cdots,d_{i,m_{i}-1}\right)$ $=\frac{\exp\left(\sum_{v=0}^{l}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}{\sum_{g=0}^{m_{i}-1}\exp\left(\sum_{v=0}^{g}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}$ (2)
@@
where $m_{i}$ is the number of response score categories for item $i$, usually 3; $x_{i}$ is the scored response to item $i$, ranging between 0 and $m_{i}-1$ (i.e., 00, $1$ and $2$, for incorrect, partially correct, and correct responses); $\theta$, $a_{i}$, $b_{i}$ have the same interpretations as in the 3PL and 2PL models; $d_{i,1}$ is the category $l$ threshold parameter. Setting $d_{i,0}=0$ and $\sum_{j=1}^{m_{i}-1}d_{i,j}=0$ resolves the indeterminacy of the model parameters.
@@
where $P_{i,l}\left(\theta\right)$ is of the form appropriate to the type of item (i.e., 3PL, 2PL or GPC), $m_{i}$ is equal to 2 for dichotomously scored items and 3 for polytomously scored items, and $u_{i,l}$ is an indicator variable defined as:
@@
This function can be viewed as a likelihood function to be maximised by the item parameters. With the estimated item parameters, $\theta$ can then be estimated (Reise & Revicki, 2014).
@@ 3.2 Test Development in Psychometrics
1. Construct and test need specification 1. (Construct and) test need specification
@@
   b. Item generation.     - Quality control.
@@
   d. Piloting of items.     - Instructions.
   e. Psychometric quality analysis.     - (Pilot) study.
4. Test construction and specification.     - Agreement analysis.
5. Implementation and testing.     - Error analysis.
@@
Test development in psychometrics concerns the process of developing and implementing a test according to psychometric principles (Irwing & Hughes, 2018). Table 1 contrasts psychometric test development (based on Irwing & Hughes (2018)) with common LLM benchmarking procedures (based on Bowman et al. (2015); Raji et al. (2021)). What sets psychometric test development apart from typical LLM benchmark development is its focus on ensuring that the test matches a well-defined construct via expert-driven item generation, rigorous pilot testing, use of factor analysis and IRT models for item and test diagnostics, establishment of scoring and normalisation standards, and testing on representative samples of intended test takers. The result of this elaborate process is a high-quality test that can assess the construct of interest for the test takers in a valid and reliable way. Many large-scale assessments, such as PISA (Programme for International Student Assessment), TIMSS and PIRLS (Progress in International Reading Literacy Study), conform to such a process.
We will use Proficiency in Grade School Mathematics (PGSM) as the construct of interest to further illustrate this process. In Step 1, the construct of interest and the test need are specified. For instance, how do we define PGSM? Is it based on a specific curriculum? What does existing literature say? Which education levels are we interested in? Is the test meant for comparison between students within a school, or between schools within a country? Such questions help us to clarify what we want to measure and how it can be measured.
@@
In Step 3, we develop test items, which is an iterative procedure involving five steps: (a) construct refinement, where we further clarify the definition of PGSM (e.g., What content domains should be included: number, algebra, probability theory? Is proficiency only about knowing, or also about applying and reasoning?); (b) generate a pool of items with domain experts; (c) review the items for obvious misfit, errors and biases; (d) pilot the items with an ideally representative sample of test takers; (e) with the responses from the pilot step, we can assess the psychometric properties of the test items with IRT and factor analysis (e.g., item discriminativeness; item difficulty; factor structure [^7] Factor structure refers to the correlational relationships between the test items expected to capture the construct of interest.). We iterate this procedure until we have a set of test items with acceptable psychometric properties. Then, in Step 4, we construct the PGSM test by specifying, for instance, which items to include (if not all), in which order, how many equivalent test versions, and what scoring instructions to use.
In Step 5, the test gets implemented to the intended test takers, followed by Step 6: another round of quality analysis. If any item displays low quality characteristics (e.g., zero or negative discriminativeness), it will be left out of the final scoring. In Step 7, responses of the test takers are scored for each item, and the resulting item-level scores form the basis for estimating proficiency scores using IRT or simpler procedures like (weighted) sums. Normalising the proficiency scores are also typical (e.g., a mean of 500 and a standard deviation of 100) to facilitate interpretations and comparisons. Finally, in Step 8, a technical manual is compiled, detailing all the results from Step 1–7, to facilitate correct re-use of the collected data, the test, as well as interpretation of test scores, among other purposes.

### v3

# PATCH! {P}sychometrics-{A}ssis{T}ed Ben{CH}marking of Large Language Models against Human Populations: A Case Study of Proficiency in 8th Grade Mathematics

@@ Abstract
Many existing benchmarks of large (multimodal) language models (LLMs) focus on measuring LLMs’ academic proficiency, often with also an interest in comparing model performance with human test takers’. While such benchmarks have proven key to the development of LLMs, they suffer from several limitations, including questionable measurement quality (e.g., Do they measure what they are supposed to in a reliable way?), lack of quality assessment on the item level (e.g., Are some items more important or difficult than others?) and unclear human population reference (e.g., To whom can the model be compared?). In response to these challenges, we propose leveraging knowledge from psychometrics—a field dedicated to the measurement of latent variables like academic proficiency—into LLM benchmarking. We make four primary contributions. First, we reflect on current LLM benchmark developments and contrast them with psychometrics-based test development. Second, we introduce PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs. PATCH addresses the aforementioned limitations. In particular, PATCH enables valid comparison between LLMs and human populations. Third, we demonstrate PATCH by measuring several LLMs’ proficiency in 8th grade mathematics against 56 human populations. We show that adopting a psychometrics-based approach yields evaluation outcomes that diverge from those based on current benchmarking practices. Fourth, we release 4 high-quality datasets to support measuring and comparing LLM proficiency in grade school mathematics and science with human populations.
PATCH! {P}sychometrics-{A}ssis{T}ed Ben{CH}marking of Large Language Models against Human Populations:
A Case Study of Proficiency in 8th Grade Mathematics
Qixiang Fang, Daniel L Oberski, Dong Nguyen, Utrecht University, Netherlands, Correspondence: q.fang@uu.nl
@@ 1 Introduction
Figure 1: PATCH: A {P}sychometrics-{A}ssis{T}ed framework for ben{CH}marking LLMs against humans.
@@
Large language models (LLMs), including their multimodal variants like vision language models, have witnessed significant advancements in recent years. These models are typically evaluated on established benchmarks that assess their performance across a diverse set of tasks such as commonsense reasoning Zellers et al. (2019); Sakaguchi et al. (2021); Chen et al. (2021), coding Chen et al. (2021); Google (2023) and academic proficiency. Academic proficiency, in particular, has become a crucial part of LLM evaluation, as evidenced by the large number of related benchmarks like MMLU, ARC, GSM8K, DROP and MATH Hendrycks et al. (2021); Clark et al. (2018); Cobbe et al. (2021); Dua et al. (2019); Hendrycks et al. (2021), as well as recent model technical reports’ increasing focus on them OpenAI (2023); Google (2023). In these benchmarks and reports, the contrast between LLM performance and human performance is often highlighted, sparking media coverage and discussions.
@@
Despite their success in advancing LLM research and shedding light on the artificial versus human intelligence debate, existing benchmarks have notable limitations. The first concern is measurement quality: Do these benchmarks measure what they are supposed to in a reliable way? Many benchmarks are created via crowd-sourced knowledge, by asking a convenience group of individuals (e.g., crowd workers, paper authors) to create new test items (e.g., GSM8K, DROP) or collecting them from (often undocumented) sources (e.g., websites, textbooks, school exams) (e.g., MATH, MMLU, ARC). Without domain expert input and rigorous testing of item quality, undesirable outcomes can occur, including a mismatch between a benchmark and its claimed measurement goal, missing information in a question, wrong answer keys, and low data annotation agreement (e.g., Nie et al., 2020; Wang et al., 2024; Chen, 2024).
@@
Second, current benchmarks do not account for differences across test items, such as item discrimination and difficulty (see Section 3.1). For example, consider three items A (easy), B (hard) and C (hard). While answering correctly to A and B would result in the same accuracy score as answering correctly to B and C, the latter (i.e., answering correctly to more difficult items) would imply higher proficiency. Furthermore, items that are too easy or too difficult (i.e., low discrimination) will fail to differentiate models (and humans) of different proficiency levels. Thus, without accounting for item differences, benchmarking results, especially model (versus human) rankings, can be misleading.
@@
Third, while many benchmarks compare LLMs against humans, the human populations under comparison remain unclear Tedeschi et al. (2023). For instance, human performance in MATH is based on the benchmark’s authors; in MMLU, crowd workers; in MATH, 6 university students. Using such convenience samples (with little information about sample characteristics), the resulting human performance cannot be generalised to other human samples or populations.
@@
To address these challenges, we propose leveraging insights from psychometrics—a field dedicated to the measurement of latent variables like academic proficiency—into LLM benchmarking practices. In particular, we draw on two research areas in psychometrics: item response theory (IRT) (Section 3.1) and test development (Section 3.2). The former enables more accurate estimation of academic proficiency on a standardised scale by taking into account both the characteristics of the test items as well as the abilities of the LLMs and individuals being assessed, compared to common practices in LLM benchmarks (e.g., using mean scores, percentages of correct responses). It can also provide diagnostic information about the quality of each test item. The latter, test development knowledge, can help to build high quality LLM benchmarks where valid comparison to specific human populations can be made (Section 3.3).
@@
Our paper makes four primary contributions. First, we reflect on current LLM benchmark development processes and contrast them with psychometrics-based test development, thereby revealing the limitations of current LLM benchmarks and the potential benefits that PATCH/psychometrics can bring to LLM benchmarking. Second, we present PATCH: a novel framework for Psychometrics-AssisTed benCHmarking of LLMs (Figure 1). PATCH is built upon IRT and test development insights from psychometrics and addresses the aforementioned limitations of existing benchmarks. Third, we demonstrate the IRT part of PATCH by testing several LLMs’ proficiency in 8th grade mathematics using the released test items and data from Trends in International Mathematics and Science Study [^1] http://timssandpirls.bc.edu/timss2015/encyclopedia/ (TIMSS) 2011. We show empirically that an IRT-based approach can lead to evaluation outcomes that diverge from those obtained through conventional benchmarking practices and that are more informative, underscoring the potential of PATCH/psychometrics to reshape the LLM benchmarking landscape. Fourth, we make our evaluation code based on the PATCH framework available [^2] https://github.com/fqixiang/patch_llm_benchmarking_with_psychometrics, along with three other mathematics and science datasets based on TIMSS 2011 and 2008333https://zenodo.org/records/12531906. See also Appendix C..
@@ 2 Related Work
We are not the first to propose leveraging psychometrics for research on LLMs and other areas in NLP. For instance, psychometric scales have been used to examine the psychological profiles of LLMs such as personality traits and motivations Huang et al. (2024); Pellert et al. (2023); Dillion et al. (2023). The text in these scales can also be used to improve encoding and prediction of personality traits Kreuter et al. (2022); Vu et al. (2020); Yang et al. (2021); Fang et al. (2023a). Psychometrics-based reliability and validity tests have also been proposed or/and used to assess the quality of NLP bias measures Du et al. (2021); van der Wal et al. (2024), text embeddings Fang et al. (2022), political stance detection Sen et al. (2020), annotations Amidei et al. (2020), user representations Fang et al. (2023b), and general social science constructs Birkenmaier et al. (2023).
The most closely related work to our paper is the use of item response theory (IRT) models in NLP for constructing more informative test datasets Lalor et al. (2016), comparison of existing evaluation datasets and instances (e.g., difficulty, discrimination) Sedoc and Ungar (2020); Vania et al. (2021); Rodriguez et al. (2021); Lalor et al. (2018); Rodriguez et al. (2022), as well as identification of difficult instances from training dynamics Lalor and Yu (2020); Lalor et al. (2019). Our work distinguishes itself from these papers in two aspects. First, we do not apply IRT to existing LLM datasets/benchmarks. Instead, we introduce a framework for benchmarking LLMs by leveraging both IRT and test development knowledge from psychometrics. The goal of this framework is to generate new, high-quality benchmarks for LLMs that warrant valid comparison with human populations. Second, we demonstrate our framework with a mathematics proficiency test validated on 56 human populations, and compare LLM performance with human performance. To the best our knowledge, we are the first to apply psychometrically validated (mathematics) proficiency tests to LLMs and make valid model versus human comparison.
@@ 3 Preliminaries
In this section, we provide background knowledge on item response theory and test development in psychometrics.
@@ 3.1 Item Response Theory
Item response theory (IRT) refers to a family of mathematical models that describe the functional relationship between responses to a test item, the test item’s characteristics (e.g., item difficulty and discrimination) and test taker’s standing on the latent construct being measured (e.g., academic proficiency) AERA et al. (2014). Unlike classical test theory and current LLM benchmarks, which focus on the total or mean score of a test, IRT models takes into account the characteristics of both the items and the individuals (and models) being assessed, offering advantages like item quality diagnostics and more accurate estimation of test takers’ proficiency. As such, IRT models have gained widespread adoption in various fields, including education, psychology, and healthcare, where trustworthy measurement and assessment are crucial.
@@
$P\left(x_{i}=1\mid\theta,a_{i},b_{i},c_{i}\right)$
$=c_{i}+\frac{1-c_{i}}{1+\exp\left(-1.7\cdot a_{i}\cdot\left(\theta-b_{i}\right)\right)}$ (1)
$\equiv P_{i,1}\left(\theta\right)$
where $x_{i}$ is the scored response to item $i$ ($1$ if correct and 00 if incorrect); $\theta$ is the proficiency of the test taker, where a higher value implies a greater probability of responding correctly; $a_{i}$ is the slope parameter of item $i$, characterising its discrimination (i.e., how well the item can tell test takers with higher $\theta$ from those with lower $\theta$) [^4] The number $1.7$ is a scaling parameter to preserve historical interpretation of parameter $a_{i}$ on the normal ogive scale Camilli (1994). Also applies to 2PL and GPC models.; $b_{i}$ is the location parameter of item $i$, characterising its difficulty; $c_{i}$ is the lower asymptote parameter of item $i$, reflecting the chances of test takers with very low proficiency selecting the correct answer (i.e., guessing). Correspondingly, the probability of an incorrect response to item $i$ is: $P_{i,0}=P\left(x_{i}=0\mid\theta_{k},a_{i},b_{i},c_{i}\right)=1-P_{i,1}\left(\theta_{k}\right)$. The 2PL model has the same form as the 3PL model (Equation 3.1), except that the $c_{i}$ parameter is fixed at zero (i.e., no guessing).
@@
$P\left(x_{i}=l\mid\theta,a_{i},b_{i},d_{i,1},\cdots,d_{i,m_{i}-1}\right)$ (2)
$=\frac{\exp\left(\sum_{v=0}^{l}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}{\sum_{g=0}^{m_{i}-1}\exp\left(\sum_{v=0}^{g}1.7\cdot a_{i}\cdot\left(\theta-b_{i}+d_{i,v}\right)\right)}$
@@
where $m_{i}$ is the number of response score categories for item $i$; $x_{i}$ is the response score of item $i$ between 0 and $m_{i}-1$ (e.g., 00, $1$ and $2$, for incorrect, partially correct, and correct responses); $\theta$, $a_{i}$, $b_{i}$ have the same interpretations as in the 3PL and 2PL models; $d_{i,1}$ is the category $l$ threshold parameter. Setting $d_{i,0}=0$ and $\sum_{j=1}^{m_{i}-1}d_{i,j}=0$ resolves the indeterminacy of the model parameters.
@@
where $P_{i,l}\left(\theta\right)$ is of the form specific to the type of item (i.e., 3PL, 2PL or GPC); $m_{i}$ equals 2 for dichotomously scored items and 3 for polytomously scored items; $u_{i,l}$ is an indicator defined as:
@@
This function can be viewed as a likelihood function to be maximised by the item parameters. With the estimated item parameters, $\theta$ can then be estimated via various algorithms Reise and Revicki (2014). In this paper, we use maximum likelihood because it gives an unbiased estimate of $\theta$.
@@ 3.2 Test Development in Psychometrics
1. Construct and test need specification. 1. (Construct and) test need specification.
@@
   b. Item generation.     – Quality control.
@@
   d. Piloting of items.     – Instructions.
   e. Psychometric quality analysis.     – (Pilot) study.
4. Test construction and specification.     – Agreement analysis.
5. Implementation and testing.     – Error analysis.
@@
Test development in psychometrics concerns the process of developing and implementing a test according to psychometric principles Irwing and Hughes (2018). Table 1 contrasts psychometric test development (based on Irwing and Hughes (2018)) with common LLM benchmarking procedures (based on Bowman et al. (2015); Raji et al. (2021)). In this section, we focus on the left panel – psychometric test development.
What sets psychometric test development apart from typical LLM benchmark development is its focus on ensuring that the test matches a well-defined construct via expert-driven item generation, rigorous pilot testing, use of factor analysis and IRT models for item and test diagnostics, establishment of scoring and normalisation standards, and testing on representative samples of intended test takers. The result of this elaborate process is a high-quality test that can assess the construct of interest for the test takers in a valid and reliable way. Many large-scale assessments, such as PISA (Programme for International Student Assessment), TIMSS and PIRLS (Progress in International Reading Literacy Study), conform to such a process.
To further illustrate this process, we propose to use the example of assessing proficiency in grade school mathematics, which is a common construct of interest in psychometric testing and LLM benchmarking. For convenience, we abbreviate this construct as PGSM.
In Step 1, the construct of interest and the test need are specified. We ask, for instance, how do we define PGSM? Is it based on a specific curriculum? What does existing literature say? Which education levels are we interested in? Is the test meant for comparison between students within a school, or between schools within a country? Such questions help us to clarify what we want to measure and how it can be measured.
@@
In Step 3, we develop test items, which is an iterative procedure involving five steps: (a) construct refinement, where we further clarify the definition of PGSM (e.g., What content domains should be included: number, algebra, and/or probability theory? Is proficiency only about knowing, or also about applying and reasoning?); (b) generate a pool of items with domain experts; (c) review the items for obvious misfit, errors and biases; (d) pilot the items with a representative sample of target test takers; (e) with the responses from the pilot step, we can assess the psychometric properties of the test items with IRT and factor analysis (e.g., item discrimination; item difficulty; factor structure [^5] Factor structure refers to the correlational relationships between test items used to measure a construct of interest.). We iterate this procedure until we have a set of test items with acceptable psychometric properties. Then, in Step 4, we construct the PGSM test by specifying, for instance, which items to include (if not all), in which order, how many equivalent test versions, and what scoring instructions to use.
In Step 5, the test gets implemented to the intended test takers, followed by Step 6: another round of quality analysis. If any item displays low quality characteristics (e.g., zero or negative discrimination), it will be left out of the final scoring. In Step 7, responses of the test takers are scored for each item, and the resulting item-level scores form the basis for estimating proficiency scores using IRT or simpler procedures like (weighted) sums. It is typical to also normalise the proficiency scores (e.g., with a mean of 500 and a standard deviation of 100) to facilitate interpretations and comparisons. Finally, in Step 8, a technical manual is compiled, detailing Step 1–7 and corresponding results, to facilitate correct re-use of the response data, the test, as well as interpretation of test scores, among other purposes.